# Survey on Large Language Model-Enhanced Reinforcement Learning: Concept, Taxonomy, and Methods

### v1

@@ VI-A Direct Decision-Maker
Reference [106] investigates the transferability of pre-trained sequence models on vision and language domains when fine-tuned on other offline tasks such as control and games. Results show that using Wikipedia-pre-trained and GPT2 language models, there are consistent performance gains in terms of reward and convergence speed in various environments. Reference [107] proposes to use pre-trained LLM as a general scaffold for task-specific model learning across various environments by adding goals along with observations as the input and converting them to sequential data. Experiments demonstrate that language modeling improves combinatorial generalization in policy learning and can substantially improve out-of-distribution performance in new tasks. To unify language reasoning with actions in a single policy, reference [108] generates textual captions interleaved with actions when training the Transformer-based policy. Results show that by using captions describing the next subgoals, the reasoning policy can consistently outperform the caption-free baseline. In-context RL is one approach to generalist RL by recalling the memory from past experiences. Reference [109] redesigns the offline in-context approach to successfully train long-sequence Transformers over entire rollouts to tackle the challenges of generalization, long-term memory, and meta-learning. To deal with situations when data is scarce and the environment is risky, reference [105] proposes Language Models for Motion Controls (LaMo) to employ pre-trained LLMs with LoRA fine-tuning method that augments the pre-trained knowledge with in-domain knowledge. The experiments indicate LaMo achieves state-of-the-art performance in sparse-reward tasks and has performance gains in data-limited scenarios. To integrate multi-modal data, e.g., vision and language, into the offline RL, reference [110] co-fine-tunes vision-language models on both robotic trajectory data and Internet-scale vision-language tasks, e.g., visual question answering. In the framework, they incorporate the actions as natural language tokens and co-fine-tune models on both vision and language datasets. Results show that such co-fine-tune methods can increase generalization performance and the chain of thought reasoning can help the agent perform multi-stage semantic reasoning and solve complex tasks.
VI-B Indirect Decision-Maker
The indirect decision-maker role is illustrated in Fig. 5 (ii). As an indirect decision-maker, LLM instructs the action selection by either generating a set of action candidates or providing a reference policy that instructs or regularizes policy learning. By instructing the action choices, LLM addresses the challenges posed by enormous action spaces and natural language in applications such as instruction-following and text-based games.
@@ VI-B1 Action Candidates
When LLM generates a set of action candidates, RL agents can further re-rank candidates based on the value function and maximize the cumulative rewards. In text-based games, action spaces are large and only a tiny fraction of action commands are accessible in any given game state. To tackle such problems, reference [111] proposes the Contextual Action Language Model (CALM) framework to generate a set of action candidates for each state in a text-based game. By training the language models on human gameplay based on the game history, the language model generates action candidates for another network to calculate the Q-values and select the optimal action. Following this work, SayCan [102] extends the application into the instruction-following agent in the robotic field. In their work, the agent integrates an LLM to generate high-level action planning and then execute low-level skills. Different from CALM, the LLM is not trained and the action candidates are generated based on prompting with the history and current observation.
VI-B2 Reference Policy
When LLM provides a reference policy using LLM, the policy updating process is modified based on the referenced policy. In human-AI collaboration, RL often converges to the equilibrium that humans may not prefer, especially in settings where human quality human data is rarely available. To solve such problems, instructRL [112] uses LLM to generate prior policy based on human instructions and uses this prior to regularizing the RL objective. Specifically, instructRL augments the policy update function with an auxiliary term $p_{\mathrm{LLM}}[\mathrm{lang(}a_{t}\mathrm{)|lang(}\tau_{t}^{i}\mathrm{),inst}]$, the probability of choosing an action based on the trajectory and instructions. Experiments show instructRL converges to policies aligned with human preferences. To solve the adaptation problem of LLM-based agents in real-time dynamic environments, Zhou et al. [113] trains a smaller-scale specialized student agent using the instruction from an LLM-based teacher agent and gradually distills the prior knowledge to the local student model, resulting in less data for training a student agent.
@@ VI-C Summarization and Outlook
As a decision-maker, LLM generates the action directly or indirectly based on observation and instructions. In direct decision-making, the sequence modeling capability of LLM is employed to improve the sample efficiency and generalization benefiting from the pre-trained world knowledge in LLM. For indirect decision-making, LLM guides the decision-making by generating action candidates or instructing the policy updating directions. In the following, challenges and future directions of the two roles are listed below:
  • Direct Decision-maker: Directly employing pre-trained large-scale LLM as the policy network demands huge computational resources and huge data for fine-tuning it in specific tasks. Future work may consider more cost-effective methods such as the Low-Rank Adaptation (LoRA) [114] to exploit the power of LLM in direct decision-making.
  • Indirect Decision-maker: Since the LLM acts as an instructor to provide reference policy, the bias and limitations are also inherited by the agent. In addition, LLM itself cannot inherently interrogate or intervene in the environment, which limits LLMs’ capabilities to intentionally aggregate information. Therefore, how to use the information gained from real-world interactions to improve the LLM itself in terms of actuality and reasoning is another important problem.
@@ VII LLM as Generator
Figure 6: LLM as a generator.
@@ VII-A1 Trajectory Rolloutor
Similar to the decision transformer [119] in offline RL, pre-trained large-scale models are used in model-based to synthesize trajectories. In 2022, Micheli et al. proposed IRIS, an agent that employs a discrete autoencoder and an autoregressive Transformer to learn the world model for Atari games [120]. With the equivalent of two hours of gameplay in the Atari 100k benchmark, the proposed method outperforms humans in 10 out of 26 games. Similarly, reference [121] applies a transformer to build a sample-efficient world model for Atari games. Utilizing such a Transformer-based world model (TWM), the RL agent can solve the long-term dependency and train a state-of-the-art policy on the Atari 100k benchmark based on generated meaningful experiences from TWM. Visual RL enables the RL agent to learn from visual observations effectively. Reference [122] propose TransDreamer, a Transformer-based model-based RL agent that leverages a Transformer for dynamics predictions in 2D and 3D visual RL tasks. Experiments show the TransDreamer agent can outperform Dreamer with long-range memory access and memory-based reasoning. Based on the development of supervised pre-training methods, Seo et al. proposes a framework that learns world model dynamics from action-free video representations [123]. The framework adds a video-based intrinsic bonus for better-guiding exploration to effectively encourage agents to learn diverse behaviors. The experimental results demonstrate their proposed method can improve the performances of vision-based RL on manipulation and locomotion tasks by transferring the pre-trained representations from unseen domains.
@@ VII-A2 Dynamics Representation Learner
Using representation learning techniques, the latent representation of the future can be learned to assist decision-making. Reference [124] introduces a visual model-based RL framework that decouples visual representation learning and dynamic learning by training an autoencoder with a vision Transformer to reconstruct pixel-given masked observations and learn the dynamics from the latent space. By such a decoupling approach, their proposed method achieves state-of-the-art performance on visual robotic tasks from Meta-world and RLBench. Utilizing the fact that language contains rich information signals and can help agents to predict the future, Lin et al. proposes Dynalang [40], where an agent that learns a multi-modal world model to predict future text and image representations and learns to act from imagined world representation. Different from other works that use language only for predicting actions, Dynalang can acquire rich language understanding by using past language and also to predict future language, videos and rewards. To solve the out-of-distribution generalization problem in visual control tasks of RL, reference [125] proposes the Language Grounded World Model (LanGWM), which focuses on learning language-grounded visual features to enhance the world model learning. To improve the generalization of the learned visual features, they masked the bounding boxes and predicted them with given language descriptions. Utilizing the expressing ability of language in higher-level concepts and global contexts, the proposed LanGWM method yields state-of-the-art results on out-of-distribution tests.
@@ VII-B Policy Interpreter
Das et al. [127] proposes a unified framework called State2Explanation (S2E), that learns a joint embedding model between state-action pairs and concept-based explanation. Based on the learned models, the explanation can help inform reward shaping during an agent’s training and provide insights to end-users at deployment. Another work  [128] first distills the policy into a decision tree, derives the decision path, and then prompts an LLM to generate a natural language explanation based on the decision path.
@@ VII-C Summarization and Outlook
  • World Model Simulator: LLMs encounter challenges in aligning their abstract knowledge with the specific requirements of different environments, leading to limitations in functional competence and grounding. This misalignment affects their effectiveness in generating trajectories and interacting with the environment. Additionally, the current model-based agents, which rely on purely observational world models, present difficulties for human adaptation. These models are typically modifiable only through observational data, which is not an effective means for humans to communicate complex intentions or adjustments. Future work may consider integrating language instructions or an adapter into the world model to achieve a flexible and adaptive world model.
  • Policy Interpreter: As policy interpreters, the quality of explanations depends on the LLM’s understanding of feature representations and implicit logic of policy. How to utilize domain knowledge or examples to improve the understanding of a complex policy is still a major issue. From the field of explainable RL, LLM may unlock its potential in different aspects. For example, from the view of feature importance, LLMs can be leveraged to interpret the correlation between observations and policy selection; from the view of the learning process and MDP, the agent’s behavior can be influenced by the MDP (such as reward $R$ and episode length $T$) or interactions with the MDP, where such influences can be revealed by asking the LLM to answer “why” and “why-not” questions with MDP context information.
@@ VIII Discussion
In previous sections, the concept ”LM-enhanced RL” and a corresponding framework have been developed and extended to the integration of multi-modal AI models such as visual-language models. Further, the different roles of LLMs in enhancing the RL paradigm have been discussed. In this section, we analyze potential applications of the LLM-enhanced RL paradigm and then discuss its future opportunities and challenges considering the untapped capabilities and limitations of LLMs, particularly in environments with multi-modal information.
VIII-A Applications of LLM-Enhanced RL
Based on the characteristics of LLM-enhanced RL, such as multi-modal information understanding and multi-task learning and generation, we believe that LLM-enhanced RL opens up a wide array of potential applications. Here we list several applications to inspire researchers.

### v3

@@ VI-A Action-Making
Several studies have further explored the application of LLMs in offline RL, demonstrating their versatility and effectiveness across various tasks and benchmarks. Reference [132] investigated the transferability of general language models on specific RL tasks. Fined-tuned on offline RL tasks (control, games), these general language models outperform Decision Transformer and reduce training time by 3-6x on D4RL benchmark [133]. Reference [129] used pre-trained LLM as a general scaffold for task-specific model learning, where goals were added along with observations as the input for LLM. Results on embodied decision-making tasks demonstrate that their proposed method outperforms others with less training data, especially when generalizing to novel tasks. In addition, they found representations in pre-trained language models can aid learning and generalization even outside of language. To unify language reasoning with actions in a single policy, reference [134] generated textual captions interleaved with actions when training the Transformer-based policy. Results show that by using captions describing the next subgoals, the reasoning policy can consistently outperform the caption-free baseline. For scenarios where data collection is costly and risky, reference [130] proposed a general framework to effectively use pre-trained LLM for offline RL. The pre-trained LLM are based on DT. To combine the pre-trained knowledge and task-related domain knowledge, they fine-tuned pre-trained LLM with Low-Rank Adaptation (LoRA) method. The architecture of Transformer is based on GPT-2 model with 12 layers and 12 attention heads. To fine-tune the LLM, they obtained the trajectory data from the D4RL dataset. Results show their method achieves state-of-the-art performance in sparse-reward tasks with limited data samples. To integrate multimodal data, e.g., vision and language, into the offline RL, reference [135] co-fine-tuned vision-language models on both robotic trajectory data and Internet-scale vision-language tasks, e.g., visual question answering. In the framework, they incorporate the actions as natural language tokens and co-fine-tune models on both vision and language datasets. Results show that such co-fine-tune methods can increase generalization performance and the chain of thought reasoning can help the agent perform multi-stage semantic reasoning and solve complex tasks.
VI-B Action-Guiding
The action-guider role is illustrated in Fig. 5 (ii). As an action-guider, LLM guides the action selection by either generating reasonable action candidates or expert actions. By instructing the action selection, LLM improves sample efficiency and exploration efficiency posed by enormous action spaces and natural language.
@@ VI-B1 Action Candidates
In environments such as text-based games, action spaces are large and only a tiny fraction of actions are accessible. Although RL agents can learn through extensive trials, they often face exploration efficiency issues, especially in multi-task settings where agents must manage various tasks simultaneously. LLMs address this challenge by generating a reduced set of action candidates based on task understanding. These candidates are likely to yield high rewards and are applicable across multiple tasks, enhancing exploration efficiency and reducing the need for ineffective exploration. Reference [136] trained a GPT-2 model to generate the candidates. To maximize long-term rewards, another neural network is used to calculate the Q-values of these candidates. When tested in 28 man-made text games from Jericho framework [137], they found the method excludes non-useful actions, speeds up the exploration and consistently achieves higher scores by more than 20%. Following this work, another study [83] proposed the SayCan framework, where instruction-following robots are integrated with an embodied LLM to understand tasks. When receiving instructions, the embodied LLM generates a high-level step-by-step plan. When acting, LLM produces action candidates based on task prompts and then the candidate with the largest critic value is executed. Being evaluated in an office kitchen with real-world robotic tasks from 101 instructions, their method can complete long-horizon, abstract, natural language instructions on a mobile manipulator.
VI-B2 Expert Actions
Traditional RL agents cannot converge to desirable equilibrium in human-AI collaboration or learn efficiently for complex tasks, due to the lack of expert demonstrations. With the understanding of human behavior and general knowledge, LLM solves the issues by producing high-quality expert actions to regularize RL agents. In human-AI collaboration, instructRL [138] used LLM to generate prior policy based on human instructions and uses this prior to regularizing the RL objective. Specifically, instructRL augments the policy update function with an auxiliary term $p_{\mathrm{LLM}}[\mathrm{lang(}a_{t}\mathrm{)|lang(}\tau_{t}^{i}\mathrm{),inst}]$, the probability of choosing an action based on the trajectory and instructions. Experiments show instructRL converges to policies aligned with human preferences. Similarly, to address the sample inefficiency issue of RL, Zhou et al. [139] included the policy difference between the student model and LLM-based teacher into RL learning loss. Their experiments on simulation platforms demonstrate that the method reduced training iterations by a factor of 1 to 9. In reference [140], LLM was leveraged to generate a high-level expert motion plan of robotics tasks, guiding RL policies to efficiently solve robotics control tasks. The high-level language plan breaks long-horizon tasks into stages to execute. Then, a single RL policy was trained across all states and stepped through the language plan. Their results show the proposed method solves long-horizon tasks from raw visual input spanning different benchmarks at success rates of over 85%, out-performing classical, language-based, and end-to-end approaches.
@@ VI-C Summarization and Outlook
Sample inefficiency and exploration inefficiency remain long-standing challenges for deep RL, particularly in environments with sparse rewards or where data collection is expensive or risky. LLM provides three ways to solve the problems. First, LLM as action-makers treat RL as a conditional sequence modeling problem. The supervised way fine-tunes pre-trained LLMs to predict future actions. Benefiting from learned prior knowledge, LLM can perform well even in out-of-distribution, sparse-reward, and long-horizon tasks. Second, LLM as action-guiders generates potential action candidates for RL. With task comprehension, these action candidates promote agents to explore potentially high task-value states, thus increasing exploration efficiency. Last, LLM generates expert actions to help RL learn from demonstrations. By incorporating demonstrations from LLM, and RL learns specific prior knowledge and improves sample efficiency.
Safety issues are another important topic when using LLMs as decision-makers in RL, particularly for costly and risky tasks [141]. For action-making, DT-based offline RL learns the optimal policy from pre-collected datasets. Recently, integrating safety constraints has been explored by some works using methods such as pessimistic estimations [142] and stationary distribution correction [143]. These methods set a constant constraint threshold before training. To dynamically adjust the threshold during deployment, a constrained decision transformer that dynamically relabels the reward based on safety-reward trade-offs was proposed [144]. For action-guiding, LLMs are viewed as instructors to provide expert actions. Ensuring the safety of actions generated by LLMs differs from that in action-making. Leveraging ideas from LLM-based agent research, we propose several potential solutions. First, LLMs can be equipped with testing modules that execute code to evaluate action safety and feasibility [145]. Second, implementing a carefully designed human-in-the-loop framework enables safety intervention through human oversight [146]. Finally, developing memory modules with reasoning mechanisms that adaptively learn safety boundaries from past experiences provides a continual learning-based approach to ensuring long-term safety [147].
In the following, we list the challenges and future directions of the two roles as below:
  • Action-making: Directly employing pre-trained large-scale LLM to generate actions demands huge computational resources and huge data for fine-tuning it in specific tasks. In the short term, future work may consider more cost-effective methods such as the Low-Rank Adaptation (LoRA) [148] to exploit the power of LLM in direct decision-making. Long-term goals involve developing innovative techniques to efficiently exploit the power of LLMs in direct decision-making, potentially creating hybrid models that combine the strengths of LLMs with more lightweight, task-specific architectures.
  • Action-guiding: Since the LLM acts as an instructor to provide expert actions, the bias and limitations are also inherited by the agent. In addition, LLM itself cannot inherently interrogate or intervene in the environment, which limits LLMs’ capabilities to intentionally aggregate information. Short-term goal is to address the inherited biases and limitations when LLMs act as instructors providing expert actions, focusing on methods to filter or correct biased information. In the long term, it is crucial to develop mechanisms for LLMs to actively interrogate and intervene in the environment, enabling them to intentionally aggregate information and improve their capabilities through real-world interactions. Therefore, how to use the information gained from real-world interactions to improve the LLM itself in terms of actuality and reasoning is another important problem.
@@ VII LLM as Generator
Figure 6: LLM as a generator. (i) World Model Simulator: LLM uses real-world data and knowledge to model dynamics, generate simulated worlds, and assist policy learning. (ii) Policy Interpreter: LLM generates interpretations of agent behavior based on state-action history and prompts, potentially leading to explainable RL.
@@ VII-A1 Trajectory Rolloutor
Similar to the decision transformer [153] in offline RL, pre-trained large-scale models were used in model-based to synthesize trajectories. In 2022, Micheli et al. proposed IRIS, an agent that employs a discrete autoencoder and an autoregressive Transformer to learn the world model for Atari games [154]. With the equivalent of two hours of gameplay in the Atari 100k benchmark, the proposed method outperforms humans in 10 out of 26 games. Similarly, reference [155] applied a transformer to build a sample-efficient world model for Atari games. Utilizing such a Transformer-based world model (TWM), the RL agent can solve the long-term dependency and train a state-of-the-art policy on the Atari 100k benchmark based on generated meaningful experiences from TWM. Visual RL enables the RL agent to learn from visual observations effectively. Reference [156] proposed TransDreamer, a Transformer-based model-based RL agent that leverages a Transformer for dynamics predictions in 2D and 3D visual RL tasks. Experiments showed the TransDreamer agent can outperform Dreamer with long-range memory access and memory-based reasoning. Based on the development of supervised pre-training methods, Seo et al. proposed a framework that learns world model dynamics from action-free video representations [157]. The framework added a video-based intrinsic bonus for better-guiding exploration to effectively encourage agents to learn diverse behaviors. The experimental results demonstrated their proposed method can improve the performances of vision-based RL on manipulation and locomotion tasks by transferring the pre-trained representations from unseen domains.
@@ VII-A2 Dynamics Representation Learner
Using representation learning techniques, the latent representation of the future can be learned to assist decision-making. Reference [114] introduced a visual model-based RL framework that decouples visual representation learning and dynamic learning by training an autoencoder with a vision Transformer to reconstruct pixel-given masked observations and learn the dynamics from the latent space. By such a decoupling approach, their proposed method achieved state-of-the-art performance on visual robotic tasks from Meta-world and RLBench. Utilizing the fact that language contains rich information signals and can help agents to predict the future, Lin et al. proposed Dynalang [46], where an agent that learns a multimodal world model to predict future text and image representations and thereby instruct the decision-making process. The algorithm is presented in Algorithm 3. In the training part, a LLM-based world model implemented with Recurrent State Space Model (RSSM) [158] computes the state representation $z_{t}$ based on the collected transitions. After that, the representation $z_{t}$ and the action $a_{t}$ are fed into the RSSM to predict the future representations $z_{t+1}$. In addition, the language $\hat{l}_{t}$ and images $\hat{x}_{t}$ are predicted to reconstruct the original state. Then, the world model is trained to learn the next representation and reconstruct observations from the representations. After updating, the world model imagined (sampled) rollouts and the policy is trained to maximize the imagined rewards. Compared to other works that use language only for predicting actions, multimodal information enables Dynalang to handle tasks that require grounded language generation, obtaining higher score than methods provided with only task descriptions. To solve the out-of-distribution generalization problem in visual control tasks of RL, reference [159] proposed the Language Grounded World Model (LanGWM), which focuses on learning language-grounded visual features to enhance the world model learning. To improve the generalization of the learned visual features, they masked the bounding boxes and predicted them with given language descriptions. Utilizing the expressing ability of language in higher-level concepts and global contexts, the proposed LanGWM method yields state-of-the-art results on out-of-distribution tests.
Algorithm 3 Training part of Dynalang [46]
0:  Rewards $r_{t}$, episode continue flag $c_{t}$, images $x_{t}$, language tokens $l_{t}$, actions $a_{t}$, model state $(h_{t},z_{t})$.
1:  while training do
2:     Draw batch of transitions $\{(r_{t},c_{t},x_{t},l_{t},a_{t})\}$ from replay buffer.
3:     Use world model to compute multimodal representations $z_{t}$, future predictions $\hat{z}_{t+1}$, and decode $\hat{x}_{t}$, $\hat{l}_{t}$, $\hat{r}_{t}$, $\hat{c}_{t}$.
4:     Update world model to minimize $\mathcal{L}_{\text{pred}}+\mathcal{L}_{\text{repr}}$.
5:     Imagine rollouts from all $z_{t}$ using $\pi$.
6:     Update actor to minimize $\mathcal{L}_{\pi}$.
7:     Update critic to minimize $\mathcal{L}_{V}$.
8:  end while
@@ VII-B Policy Interpreter
Das et al. [161] proposed a unified framework called State2Explanation (S2E), that learns a joint embedding model between state-action pairs and concept-based explanation. Based on the learned models, the explanation can help inform reward shaping during an agent’s training and provide insights to end-users at deployment. Another work  [162] first distilled the policy into a decision tree, derives the decision path, and then prompts an LLM to generate a natural language explanation based on the decision path. Additionally, Lu et al. [163] introduced a framework that decomposes the overall reward into multiple sub-rewards based on specific object properties, defines actions as high-level motion primitives executed at precise 3D positions to simplify decision-making, and integrates LLMs to enable interactive and flexible querying of explanations.
@@ VII-C Summarization and Outlook
  • World Model Simulator: LLMs encounter challenges in aligning their abstract knowledge with the specific requirements of different environments, leading to limitations in functional competence and grounding. This misalignment affects their effectiveness in generating trajectories and interacting with the environment. Additionally, the current model-based agents, which rely on purely observational world models, present difficulties for human adaptation. These models are typically modifiable only through observational data, which is not an effective means for humans to communicate complex intentions or adjustments. Furthermore, while LLMs hold promise for enhancing multi-task learning by generating trajectories and dynamic representations applicable to multiple tasks, there remains a scarcity of research exploring this potential.Looking ahead, in the short term, researchers might prioritize improving LLMs’ alignment with specific environment requirements. For the long term, a promising direction would be to integrate language instructions or an adapter into the world model. This approach could lead to a more flexible and adaptive world model that can better accommodate human intentions and adjustments, as well as support more effective multi-task learning through enhanced trajectory generation and dynamic representations.
  • Policy Interpreter: As policy interpreters, the quality of explanations depends on the LLM’s understanding of feature representations and implicit logic of policy. How to utilize domain knowledge or examples to improve the understanding of a complex policy is still a major issue. In the short term, researchers could focus on enhancing LLMs’ ability to interpret the correlation between observations and policy selection, providing insights into the decision-making process of RL agents. This could involve developing techniques to better leverage domain knowledge and examples for improving LLMs’ understanding of complex policies. In the long-term, the field could explore advanced applications of LLMs in explainable RL, such as analyzing the learning process and MDP to reveal influences on agent behavior. Researchers could develop sophisticated prompting techniques for LLMs to answer nuanced “why” and “why-not” questions with MDP context, providing deeper explanations of agent decision-making.
@@ VIII Discussion
In previous sections, we introduced the concept “LLM-enhanced RL” and developed a corresponding framework, which we extended to the integration of multimodal AI models such as visual-language models. We then discussed the different LLM-enhanced RL approaches. This section provides a comprehensive analysis of the LLM-enhanced RL approach. We begin with a comparative analysis of the different LLM roles, highlighting their advantages and limitations. Following this, we explore potential real-world applications of the LLM-enhanced RL paradigm. Finally, we discuss future opportunities and challenges, taking into account both the untapped capabilities and inherent limitations of LLMs, with a particular focus on their application in multimodal information environments.
VIII-A Comparison of Different LLM-Enhanced RL Approaches
This section provides a comparative analysis of the four LLM-enhanced RL approaches, helping researchers understand the strengths and limitations of each approach.
  • Information Processor: As an information processor, LLM excels in handling complex, multimodal inputs, particularly in translating natural language instructions or environment information into a format more readily usable by RL agents. This role significantly enhances the agent’s ability to understand and interact with complex environments. However, it faces challenges in computational efficiency and may struggle with highly specialized domain knowledge not covered in its pre-training.
  • Reward Designer: LLMs serving as reward designers offer a more intuitive and flexible approach to defining reward functions, especially in complex or sparse-reward environments. This role can significantly improve the alignment of RL objectives with human intentions. The main limitation lies in ensuring that generated rewards accurately reflect task-specific nuances and long-term goals, particularly in highly specialized domains.
  • Decision-Maker: As decision-makers, LLMs can either directly generate actions or guide action selection, leveraging their vast knowledge base to improve sample efficiency and exploration in RL. This role is particularly effective in tasks requiring complex reasoning or long-term planning. However, it may face challenges in real-time decision-making scenarios due to computational overhead and may inherit biases present in the LLM’s training data.
  • Generator: In the generator role, LLMs can simulate complex environments for model-based RL and provide interpretable explanations of RL policies. This capability is invaluable for improving sample efficiency and making RL more transparent and understandable. The main challenges include aligning generated simulations with real-world dynamics and ensuring the relevance and accuracy of policy explanations.
VIII-B Applications of LLM-Enhanced RL
Based on the characteristics of LLM-enhanced RL, such as multimodal information understanding and multi-task learning and generation, we believe that LLM-enhanced RL opens up a wide array of potential applications. Here we list several applications to inspire researchers.